Frank made 144 pounds of coffee for a total cost of $666.00. Coffee A cost $5.45 per pound, and coffee B cost $4.10 per pound. How much of each coffee did Frank use?
Found 2 solutions by Boreal, jorel555:
Answer by Boreal(15235)

A+B=144
5.45A+4.10B=666
Multiply the top by -4.10
-4.10A-4.10B=-590.4
5.45A+4.10B=666
1.35A=75.6
A=56 pounds@$5.45/lb=$305.20
B=88 pounds@$4.10/lb=$360.80
That is $666.00

Answer by jorel555(1290)

Let n be the amount of coffee A. Then:
545n+410(144-n)=66600
545n+59040-410n=66600
135n=7560
n=56 lbs of coffee A, 88 lbs of coffee B!!!!!!!!!!!!